OpenStudy (anonymous):

Simplify: 3y^2-4y-7/ 9y^2 divided by y^2-6y-7/3y^2-14y-49

OpenStudy (anonymous):

just a pro tip: www.wolframalpha.com Calculating atm will return answer asap.

OpenStudy (anonymous):

i need to know how to solve it. like steps

OpenStudy (anonymous):

it doesnt understand the ? on wolfram because I dont know how to type 3y^2-4y-7 over 9y^2-49

OpenStudy (anonymous):

((3y^2-4y-7)/(9y^2))/((y^2-6y-7)/(3y^2-14y-49)) like this.

OpenStudy (anonymous):

factorise everything

OpenStudy (anonymous):

3y^2 -4y -7

OpenStudy (anonymous):

two numbers multiply to give -21 and add to give -4

OpenStudy (anonymous):

-7 and +3

OpenStudy (anonymous):

3y^2 +3y-7y -7 3y(y+1) -7(y+1) = (y+1)(3y-7)

OpenStudy (anonymous):

etc, do for the rest

OpenStudy (anonymous):

y^2-6y-7 = (y-7)(y+1)

OpenStudy (anonymous):

3y^2 -14y -49

OpenStudy (anonymous):

factor that , you can do that llook for two numbers that multiply to give (3 x -49 ) and add to give -14

OpenStudy (anonymous):

I got this...

OpenStudy (anonymous):

:|

OpenStudy (anonymous):

yeh, but you did nothing lols, all u did was rewrite it you are meant to factor everything, then stuff will cancel

OpenStudy (anonymous):

3y^2 -14y -49 = (y-7)(3y+7)

OpenStudy (anonymous):

then lots cancels

OpenStudy (anonymous):

\[= \frac{ (y+1)(3y-7) }{9y^2} \times \frac{(y-7)(3y+7) }{(y-7)(y+1)}\]

OpenStudy (anonymous):

\[= \frac{ 9y^2 -49}{9y^2}\]

OpenStudy (anonymous):

\[= 1 - \frac{49}{9y^2}\] if you really want